### v1

# Distributionally Robust Safe Sample Screening

Abstract
In this study, we propose a machine learning method called Distributionally Robust Safe Sample Screening (DRSSS). DRSSS aims to identify unnecessary training samples, even when the distribution of the training samples changes in the future. To achieve this, we effectively combine the distributionally robust (DR) paradigm, which aims to enhance model robustness against variations in data distribution, with the safe sample screening (SSS), which identifies unnecessary training samples prior to model training. Since we need to consider an infinite number of scenarios regarding changes in the distribution, we applied SSS because it does not require model training after the change of the distribution. In this paper, we employed the covariate shift framework to represent the distribution of training samples and reformulated the DR covariate-shift problem as a weighted empirical risk minimization problem, where the weights are subject to uncertainty within a predetermined range. By extending the existing SSS technique to accommodate this weight uncertainty, the DRSSS method is capable of reliably identifying unnecessary samples under any future distribution within a specified range. We provide a theoretical guarantee for the DRSSS method and validate its performance through numerical experiments on both synthetic and real-world datasets.
1 Introduction
In practical applications of machine learning, there are cases where the data distribution differs between the training and deployment phases. If the data distribution at deployment is unknown, it becomes necessary to robustly handle a certain range of distributional changes, which is referred to as distributionally robust (DR) learning (see, e.g., [1]). In standard DR learning settings, the focus is on how robustly a model fixed at the training phase behaves when the data distribution changes at deployment. However, in practice, there are cases where fine-tuning or retraining of the model is possible at deployment. In such settings, it is preferable to minimize the data storage costs and the retraining costs at deployment. This issue becomes more critical in scenarios such as continual learning (see, e.g., [2]), where the model needs to be continually updated in response to environmental changes. In this study, we propose a method called distributionally robust safe sample screening (DRSSS) method that allows for removing a set of training samples in advance while ensuring the optimality of the retrained model in the deployment phase where the data distribution changes within a certain range.
The basic idea of the proposed DRSSS method is to extend an optimization technique called safe sample screening (SSS) to the DR learning setting. SSS is an approach that allows the removal of some training instances in advance for learning sparse models such as support vector machines (SVM) [3]. This approach is termed safe because it guarantees that the solution to the optimization problem, even after removing those training samples, will be exactly the same as the solution obtained with all training samples. Various SSS methods have been developed for different models, demonstrating that they can make the learning process of sparse models more efficient. However, our goal in this study is not merely to improve efficiency, but to safely identify unnecessary samples that are robust to distributional changes. We develop a method that ensures optimality after data reduction, regardless of how the data distribution varies within a certain range, by extending the SSS technology to the DR learning setting. While traditional SSS only needs to guarantee the optimality for a single optimal solution, DRSSS is challenging because it requires guaranteeing optimality for an entire class (i.e., an infinite number of optimization problems) across a certain range of distribution changes.
In this study, we consider the covariate shift setting as the distributional change at deployment phases. Covariate shift is a scenario where the distribution of input data changes between the training phase and the test (deployment) phase [4, 5]. If the input data distribution during the deployment phase is known, the problem of covariate shift can be addressed as a weighted optimization problem by assigning weights to each training sample (formal definition we employed is presented in Section 2.1). The weight of each sample can be determined according to the ratio of the density in the test (deployment) distribution to that in the training distribution. In the DR learning setting of this study, we assume that the test input distribution is unknown but exists within a certain range from the training input distribution. This means that the problem of DR covariate shift can be formulated as an optimization problem with DR weighting, where the weights of each sample vary within a certain range. Our proposed DRSSS method enables SSS for this DR weighting optimization problem (Section 3). Figure 1 illustrates the proposed DRSSS method.
Figure 1: Schematic illustration of the proposed DRSSS method in a toy binary classification problem with a sample-sparse classifier such as SVM. (A) Suppose that we have a dataset of binary classification (classes described by colors), and we compute the proposed DRSSS in the initial training phase. Then, it can identify a set of samples that do not influence the optimal solutions in the deployment phase, regardless of any changes within a specified range in the input distribution (points with thin colors). (B) This means that, for any input distribution of the deployment phase within a specified range (depicted by the differing sizes of points), the set of samples screened out by the DRSSS method is always identified as no influence (screened samples in a specific deployment is always a superset of that in DRSSS).
Related works
The DR setting has been explored in various machine learning problems, aiming to enhance model robustness against data distribution variations. A DR learning problem is typically formulated as a worst-case optimization problem since the goal is to ensure model performance under the worst-case data distribution within a specified range. Consequently, a variety of optimization techniques tailored to DR learning have been investigated within both the machine learning and optimization communities [6, 7, 1]. The proposed DRSSS method is one such DR learning method, focusing specifically on the problem of sample reduction. The ability to identify unnecessary samples is of practical significance. For example, in the context of continual learning (see, e.g., [2]), it is crucial to effectively manage data by selectively retaining and discarding samples, especially in anticipation of changes in future data distributions. Incorrect deletion of essential data can lead to catastrophic forgetting [8], a phenomenon where a model, after being trained on new data, quickly loses information previously learned from older datasets. The proposed DRSSS method addresses this challenge by identifying samples that, regardless of future data distribution shifts, will not have any influence on all possible newly trained models in the future.
Safe screening (SS) refers to optimization techniques in sparse learning that identify and exclude unnecessary samples or features from the learning process, thereby reducing computational cost without altering the final trained model. Initially, safe feature screening was introduced by [9] for the Lasso, followed by safe sample screening by [3] for the SVM. Among various SS methods developed so far, the most commonly used is based on the duality gap [10, 11, 12], which is also used for the proposed DRSSS method. Over the past decade, SS has seen diverse developments, including methodological improvements and expanded application scopes [12, 13, 14, 15, 16, 17, 18, 19]. Unlike other SSS studies that primarily focused on reducing computational costs, this study adopts SSS for a different purpose. We employ SSS to identify unnecessary samples from an infinitely large number of learning problems, or more specifically, across infinitely large numbers of scenarios where data distribution varies within a specified range. To our knowledge, no existing studies have utilized SS within the DR learning framework.
Contributions
Our contributions are summarized as follows. First, by effectively combining distributionally robust (DR) learning and safe sample screening (SSS), we propose a novel method called the DRSSS method that can identify samples that do not affect the optimal solutions in the deployment phase, regardless of how the distribution changes within a specified range. Although the DRSSS method is developed for convex empirical risk minimization problems, we have extended the proposed method through kernelization to be applicable to complex nonlinear models. Using the neural tangent kernel, this extension enables the application of DRSSS to practical deep learning models. Finally, through numerical experiments, we verify the effectiveness of the proposed DRSSS method.
The code for the experiment is available as supplementary information.
Notations
Notations used in this paper are described in Table 1.
Table 1: Notations used in the paper. $\mathbb{R}$: all real numbers, $\mathbb{N}$: all positive integers, $n,n^{\prime},p\in\mathbb{N}$, $f:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$: convex function, $M\in\mathbb{R}^{n\times n^{\prime}}$: matrix, $\bm{v}\in\mathbb{R}^{n}$: vector.
$m_{ij}\in\mathbb{R}$ (small case of matrix variable) the element at the $i^{\mathrm{th}}$ row and the $j^{\mathrm{th}}$ column of $M$
$v_{i}\in\mathbb{R}$ (nonbold font of vector variable) the $i^{\mathrm{th}}$ element of $\bm{v}$
$M_{i:}\in\mathbb{R}^{1\times n^{\prime}}$ the $i^{\mathrm{th}}$ row of $M$
$M_{:j}\in\mathbb{R}^{n\times 1}$ the $j^{\mathrm{th}}$ column of $M$
$[n]$$\{1,2,\dots,n\}$
$\mathbb{R}_{\geq 0}$ all nonnegative real numbers
$\otimes$ elementwise product
diagonal matrix: $\mathrm{diag}(\bm{v})\in\mathbb{R}^{n\times n}$ $(\mathrm{diag}(\bm{v}))_{ii}=v_{i}$ and $(\mathrm{diag}(\bm{v}))_{ij}=0$ ($i\neq j$)
$\bm{v}{\raisebox{0.80002pt}{$\times$}\Box}M\in\mathbb{R}^{n\times n^{\prime}}$ $\mathrm{diag}(\bm{v})M$
$\bm{0}_{n}\in\mathbb{R}^{n}$ $[0,0,\dots,0]^{\top}$ (vector of size $n$)
$\bm{1}_{n}\in\mathbb{R}^{n}$ $[1,1,\dots,1]^{\top}$ (vector of size $n$)
$p$-norm: $\|\bm{v}\|_{p}\in\mathbb{R}_{\geq 0}$ $(\sum_{i=1}^{n}v_{i}^{p})^{1/p}$
subgradient: $\partial f(\bm{v})\subseteq\mathbb{R}^{n}$$\{\bm{g}\in\mathbb{R}^{n}\mid\forall\bm{v}^{\prime}\in\mathbb{R}^{n}:~{}f(\bm{v}^{\prime})-f(\bm{v})\geq\bm{g}^{\top}(\bm{v}^{\prime}-\bm{v})\}$
${\cal Z}[f]\subseteq\mathbb{R}^{n}$$\{\bm{v}^{\prime}\in\mathbb{R}^{n}\mid\partial f(\bm{v}^{\prime})=\{\bm{0}_{n}\}\}$
convex conjugate: $f^{*}(\bm{v})\in\mathbb{R}\cup\{+\infty\}$ $\sup_{\bm{v}^{\prime}\in\mathbb{R}^{n}}(\bm{v}^{\top}\bm{v}^{\prime}-f(\bm{v}^{\prime}))$
$f$ is $\kappa$-strongly convex” ($\kappa>0$) $f(\bm{v})-(\kappa/2)\|\bm{v}\|_{2}^{2}$ is convex with respect to $\bm{v}$
$f$ is $\mu$-smooth” ($\mu>0$) $\|f(\bm{v})-f(\bm{v}^{\prime})\|_{2}\leq\mu\|\bm{v}-\bm{v}^{\prime}\|_{2}$ for any $\bm{v},\bm{v}^{\prime}\in\mathbb{R}^{n}$
2 Preliminaries
2.1 Weighted regularized empirical risk minimization
We first present the class of machine learning problem that the covariate shift setting is acceptable and SSS is applicable. As noted in Section 1, a machine learning model applicable to the covariate shift setting can be implemented by assigning weights to each training sample. Thus we employ the class of the machine learning model called the weighted regularized empirical risk minimization (weighted RERM), for linear predictions. We note that this “linear prediction” includes the kernel method, which is discussed in Section 4.2.
Definition 2.1.
Given $n$ training samples of $d$-dimensional input variables, scalar output variables and scalar sample weights, denoted by $X\in\mathbb{R}^{n\times d}$, $\bm{y}\in\mathbb{R}^{n}$ and $\bm{w}\in\mathbb{R}_{\geq 0}^{n}$, respectively, the training computation of weighted RERM for linear prediction is formulated as follows:
$\bm{\beta}^{*(\bm{w})}:=\mathop{\rm argmin}\limits_{\bm{\beta}\in\mathbb{R}^{d}}P_{\bm{w}}(\bm{\beta}),\quad\text{where}\quad P_{\bm{w}}(\bm{\beta}):=\sum_{i=1}^{n}w_{i}\ell_{y_{i}}(\check{X}_{i:}\bm{\beta})+\rho(\bm{\beta}).$ (1)
Here, $\ell_{y}:\mathbb{R}\to\mathbb{R}$ is a convex loss function [^1] For $\ell_{y}(t)$, we assume that only $t$ is a variable of the function ($y$ is assumed to be a constant) when we take its subgradient or convex conjugate., $\rho:\mathbb{R}^{d}\to\mathbb{R}$ is a convex regularization function, and $\check{X}\in\mathbb{R}^{n\times d}$ is a matrix calculated from $X$ and $\bm{y}$ and determined depending on $\ell_{y}$. In this paper, unless otherwise noted, we consider binary classifications ($\bm{y}\in\{-1,+1\}^{n}$) with $\check{X}:=\bm{y}{\raisebox{0.80002pt}{$\times$}\Box}X$. For regressions ($\bm{y}\in\mathbb{R}^{n}$) we usually set $\check{X}:=X$.
Here, $\bm{\beta}^{*(\bm{w})}$ is the linear prediction coefficients: given a sample $\hat{\bm{x}}\in\mathbb{R}^{d}$, its prediction is assumed to be done by $\hat{\bm{x}}^{\top}\bm{\beta}^{*(\bm{w})}$. Since $\ell_{y}$ and $\rho$ are convex, we can easily confirm that $P_{\bm{w}}(\bm{\beta})$ is convex with respect to $\bm{\beta}$.
Remark 2.2.
We add that, we adopt the formulation $X_{:d}=\bm{1}_{n}$ so that $\beta^{*(\bm{w})}_{d}$ (the last element) represents the common coefficient for any sample (called the intercept).
Applying Fenchel’s duality theorem (Appendix B), we have the following dual problem of (1):
$\bm{\alpha}^{*(\bm{w})}:=\mathop{\rm argmax}\limits_{\bm{\alpha}\in\mathbb{R}^{n}}D_{\bm{w}}(\bm{\alpha}),\quad\text{where}\quad D_{\bm{w}}(\bm{\alpha}):=-\sum_{i=1}^{n}w_{i}\ell^{*}_{y_{i}}(-\alpha_{i})-\rho^{*}((\bm{w}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})^{\top}\bm{\alpha}).$ (2)
The relationship between the original problem (1) (called the primal problem) and the dual problem (2) are described as follows:
$P_{\bm{w}}(\bm{\beta}^{*(\bm{w})})=D_{\bm{w}}(\bm{\alpha}^{*(\bm{w})}),$ (3)
$\bm{\beta}^{*(\bm{w})}\in\partial\rho^{*}((\bm{w}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})^{\top}\bm{\alpha}^{*(\bm{w})}),$ (4)
$\forall i\in[n]:\quad-\alpha^{*(\bm{w})}_{i}\in\partial\ell_{y_{i}}(\check{X}_{i:}\bm{\beta}^{*(\bm{w})}).$ (5)
Remark 2.3.
We should notice that, depending on $\bm{\alpha}$, $D_{\bm{w}}(\alpha)$ may be infinite and therefore we may not use such $\bm{\alpha}$ when computing it. This is because the convex conjugate $f^{*}$ (Table 1) may be infinite even $f$ is finite. In this setup, since we do not use $\rho$ such that $\rho^{*}$ may be infinite (discussed in Corollary A.3 in Appendix A), we have to care about only the condition when $\ell^{*}_{y_{i}}(-\alpha_{i})$ becomes infinite. In the examples of loss functions in Table 2 we present conditions for $\ell^{*}_{y_{i}}(-\alpha_{i})<+\infty$.
2.2 Sample-sparse loss functions
In weighted RERM, we call that a loss function $\ell_{y}$ induces sample-sparsity if elements in $\bm{\alpha}^{*(\bm{w})}$ are easy to become zero. As implied from (4), if $\alpha^{*(\bm{w})}_{i}=0$, the $i^{\mathrm{th}}$ sample (i.e., the $i^{\mathrm{th}}$ row of $\check{X}$) is unnecessary to represent the primal model parameter $\bm{\beta}^{*(\bm{w})}$ and therefore we can remove them. Due to (5), the sample-sparsity can be achieved by $\ell_{y}$ such that ${\cal Z}[\ell_{y}]:=\{t\in\mathbb{R}\mid\partial\ell_{y}(t)=\{0\}\}$ is not a point but an interval.
In case of binary classification problems ($y\in\{-1,+1\}$), the hinge loss $\ell_{y}(t)=\max\{0,1-t\}$ is a famous sample-sparse loss function. In fact, ${\cal Z}[\ell_{y}]=(1,+\infty)$ for the hinge loss. Other examples are presented in Table 2.
Table 2: Examples of sample-sparse loss functions. Calculation results used later are also presented.
$\ell_{y}(t)$ $y\in$ $\check{X}\phantom{\Bigr{]}}$ $\ell^{*}_{y}(t)$ $\ell^{*}_{y_{i}}(-\alpha_{i})<+\infty$ if ${\cal Z}[\ell_{y}]$
$\max\{0,1-t\}$$\{-1,+1\}$ $\bm{y}{\raisebox{0.80002pt}{$\times$}\Box}X$ $t$ ($-1\leq t\leq 0$) $0\leq\alpha_{i}\leq 1$ $(1,+\infty)$
   (Hinge loss) +$\infty$+ ∞ (otherwise)
$(\max\{0,1-t\})^{2}$$\{-1,+1\}$ $\bm{y}{\raisebox{0.80002pt}{$\times$}\Box}X$ $(t^{2}+4t)/4$ ($t\leq 0$) $\alpha_{i}\geq 0$ $[1,+\infty)$
   (Squared hinge loss) +$\infty$+ ∞ (otherwise)
[…]
   ($\varepsilon$-insensitive loss; $t(y+\varepsilon)$ ($0\leq t\leq 1$)
   $\varepsilon>0$: hyperparameter) +$\infty$+ ∞ (otherwise)
[…]
   (squared $\varepsilon$-insensitive loss;
   $\varepsilon>0$: hyperparameter)
3 Distributionally robust safe sample screening
In this section we show the condition of unnecessary samples identified by DRSSS, which we call the DRSSS rule, for weighted RERM with two steps. First, in Sections 3.1 we show SSS rules for weighted RERM but not DR setup. Here, “DR setup” means that the weight is assumed to be changed in a predetermined range. To do this, we extended existing SSS rules in [11, 12]. Then we derive DRSSS rules in Section 3.2.
3.1 (Non-DR) safe sample screening
We consider identifying training samples that do not affect the learned primal model parameter $\bm{\beta}^{*(\bm{w})}$. As discussed in Section 2.2, such samples can be identified as $i\in[n]$ such that $\alpha^{*(\bm{w})}_{i}=0$. However, since computing $\bm{\alpha}^{*(\bm{w})}$ is as costly as $\bm{\beta}^{*(\bm{w})}$, we would like to judge whether $\alpha^{*(\bm{w})}_{i}=0$ or not without computing exact $\bm{\alpha}^{*(\bm{w})}$. To solve the problem, the SSS first considers identifying the possible range ${\cal B}^{*(\bm{w})}\subset\mathbb{R}^{d}$ such that $\bm{\beta}^{*(\bm{w})}\in{\cal B}^{*(\bm{w})}$ is assured. Then, with ${\cal B}^{*(\bm{w})}$ and (5), we can conclude that the $i^{\mathrm{th}}$ training sample do not affect the training result $\bm{\beta}^{*(\bm{w})}$ if $\bigcup_{\bm{\beta}\in{\cal B}^{*(\bm{w})}}\partial\ell_{y_{i}}(\check{X}_{i:}\bm{\beta})=\{0\}$.
First we show how to compute ${\cal B}^{*(\bm{w})}$. In this paper we adopt the computation methods that is available when the regularization function $\rho$ in $P_{\bm{w}}$ (and also $P_{\bm{w}}$ itself) of (1) are strongly convex.
Lemma 3.1.
Suppose that $\rho$ in $P_{\bm{w}}$ (and also $P_{\bm{w}}$ itself) of (1) are $\kappa$-strongly convex. Then, for any $\hat{\bm{\beta}}\in\mathbb{R}^{d}$ and $\hat{\bm{\alpha}}\in\mathbb{R}^{n}$, we can assure that the following ${\cal B}^{*(\bm{w})}\subset\mathbb{R}^{d}$ must satisfy $\bm{\beta}^{*(\bm{w})}\in{\cal B}^{*(\bm{w})}$:
${\cal B}^{*(\bm{w})}:=\{\bm{\beta}\in\mathbb{R}^{d}\mid\|\bm{\beta}-\hat{\bm{\beta}}\|_{2}\leq r(\bm{w},\kappa,\hat{\bm{\beta}},\hat{\bm{\alpha}})\},$
$\text{where}\quad r(\bm{w},\kappa,\hat{\bm{\beta}},\hat{\bm{\alpha}}):=\sqrt{(2/\kappa)[P_{\bm{w}}(\hat{\bm{\beta}})-D_{\bm{w}}(\hat{\bm{\alpha}})]}.$
The proof is done by extending [11, 12] to our setup; see Appendix C.1. The quantity $P_{\bm{w}}(\hat{\bm{\beta}})-D_{\bm{w}}(\hat{\bm{\alpha}})$ is known as the duality gap, which must be nonnegative due to (3). Then we obtain the following SSS rule, which is referred to as the gap safe screening rule [11]:
Lemma 3.2.
Under the same assumptions as Lemma 3.1, $\alpha_{i}^{*(\bm{w})}=0$ is assured (i.e., the $i^{\mathrm{th}}$ training sample does not affect the training result $\bm{\beta}^{*(\bm{w})}$) if there exists $\hat{\bm{\beta}}\in\mathbb{R}^{d}$ and $\hat{\bm{\alpha}}\in\mathbb{R}^{n}$ such that
$[\check{X}_{i:}\hat{\bm{\beta}}-\|\check{X}_{i:}\|_{2}r(\bm{w},\kappa,\hat{\bm{\beta}},\hat{\bm{\alpha}}),~{}\check{X}_{i:}\hat{\bm{\beta}}+\|\check{X}_{i:}\|_{2}r(\bm{w},\kappa,\hat{\bm{\beta}},\hat{\bm{\alpha}})]\subseteq{\cal Z}[\ell_{y_{i}}].$
The proof is presented in Appendix C.2.
Remark 3.3.
When we employ the intercept in the model (Remark 2.2), we often define that the regularization function $\rho(\bm{\beta})$ do not regularize $\beta_{d}$ (the model parameter for the intercept), that is, $\rho(\bm{\beta})$ is constant with respect to $\beta_{d}$. However, to apply Lemmas 3.1 and 3.2, we need to define that $\rho(\bm{\beta})$ does regularize $\beta_{d}$, because $\rho$ itself is required to be $\kappa$-strongly convex.
3.2 Extension of SSS to DR setup
With the conditions when samples are screened out proved in Section 3.1, we then define the problem setup of DRSSS.
Definition 3.4 (Distributionally robust safe sample screening (DRSSS)).
Given $X\in\mathbb{R}^{n\times d}$, $\bm{y}\in\mathbb{R}^{n}$ and ${\cal W}\subset\mathbb{R}_{\geq 0}^{n}$, the DRSSS for ${\cal W}$ in weighted RERM is defined as finding unnecessary samples for any $\bm{w}\in{\cal W}$, or more specifically, finding $i\in[n]$ that satisfies Lemma 3.2 for any $\bm{w}\in{\cal W}$.
To do this, if Lemma 3.2 is applicable, we have only to take the following procedure:
Theorem 3.5.
If $\rho$ in $P_{\bm{w}}$ (1) is $\kappa$-strongly convex, DRSSS in Definition 3.4 is achieved as follows:
  1. 1.
    Choose $\tilde{\bm{w}}\in{\cal W}$.
  2. 2.
    Compute (1)&(5) or (2)&(4) for $\tilde{\bm{w}}$ to learn $\bm{\beta}^{*(\tilde{\bm{w}})}$ and $\bm{\alpha}^{*(\tilde{\bm{w}})}$.
  3. 3.
    Identify $i\in[n]$ that satisfies Lemma 3.2 with $\hat{\bm{\beta}}\leftarrow\bm{\beta}^{*(\tilde{\bm{w}})}$ and $\hat{\bm{\alpha}}\leftarrow\bm{\alpha}^{*(\tilde{\bm{w}})}$ and for any $\bm{w}\in{\cal W}$. Such $i$ assures $\alpha_{i}^{*(\bm{w})}=0$ $\forall\bm{w}\in{\cal W}$.
Here, the last procedure can be computed as follows:
$[\check{X}_{i:}\bm{\beta}^{*(\tilde{\bm{w}})}-\|\check{X}_{i:}\|_{2}R,~{}\check{X}_{i:}\bm{\beta}^{*(\tilde{\bm{w}})}+\|\check{X}_{i:}\|_{2}R]\subseteq{\cal Z}[\ell_{y_{i}}],$ (6)
$\text{where}\quad R:=\max_{\bm{w}\in{\cal W}}r(\bm{w},\kappa,\bm{\beta}^{*(\tilde{\bm{w}})},\bm{\alpha}^{*(\tilde{\bm{w}})})=\sqrt{(2/\kappa)\max_{\bm{w}\in{\cal W}}[P_{\bm{w}}(\bm{\beta}^{*(\tilde{\bm{w}})})-D_{\bm{w}}(\bm{\alpha}^{*(\tilde{\bm{w}})})]}.$ (7)
In Theorem 3.5, the key point is the maximization in (7), which is a maximization of convex function and therefore is difficult in general. In fact,
$P_{\bm{w}}(\bm{\beta}^{*(\tilde{\bm{w}})})-D_{\bm{w}}(\bm{\alpha}^{*(\tilde{\bm{w}})})$
$=\underbrace{\sum_{i=1}^{n}w_{i}[\ell_{y_{i}}(\check{X}_{i:}\bm{\beta}^{*(\tilde{\bm{w}})})+\ell^{*}_{y_{i}}(-\alpha^{*(\tilde{\bm{w}})}_{i})]}_{\text{linear w.r.t.}~{}\bm{w}}+\rho(\bm{\beta}^{*(\tilde{\bm{w}})})+\underbrace{\rho^{*}((\bm{w}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})^{\top}\bm{\alpha}^{*(\tilde{\bm{w}})})}_{\text{convex w.r.t.}~{}\bm{w}},$ (8)
is convex with respect to $\bm{w}$. (Note that $\rho^{*}$ is assured to be convex.) However, depending on the choice of the regularization function $\rho$ and the range of weight changes ${\cal W}$, it may be explicitly computed. So, in Section 4 we show specific calculations for some typical setups.
4 Specific computations of DRSSS for L2-regularization
In this section we discuss specific computations of DRSSS in Theorem 3.5, which depends on the choice of the loss function $\ell_{y}$, the regularization function $\rho$, and the distribution changes (i.e., range of weight changes ${\cal W}$). Especially, we present how to compute the quantities to conduct DRSSS: computation of (6) and maximization of (8) in the theorem.
First we can see that the choice of the loss function does not affect the required computations for DRSSS so much, since the terms including the loss function $\ell_{y}$ and its convex conjugate $\ell^{*}_{y}$ in (6) and (8) are calculated only by known quantities $\check{X}$, $\bm{\beta}^{*(\tilde{\bm{w}})}$ and $\bm{\alpha}^{*(\tilde{\bm{w}})}$. So, after determining $\ell_{y}$, we have only to apply functions described in Table 2.
On the other hand, the choice of the regularization function and the range of weight changes largely affect the difficulty of the computations. So we mainly discuss the use of L2-regularization and show the algorithm to compute them. Note that, it is not obvious to extend DRSSS for L2-regularization to other regularization functions; discussions are presented in Appendix D.1.
4.1 DRSSS for L2-regularization and weight changes in L2-norm
As stated in Lemma 3.1, DRSSS requires the regularization function to be $\kappa$-strongly convex. So the L2-regularization $\rho(\bm{\beta}):=\frac{\lambda}{2}\|\bm{\beta}\|_{2}^{2}$ ($\lambda>0$) can be applicable for DRSSS. In addition, we assume that the range of weight changes ${\cal W}$ is defined by L2-norm ${\cal W}:=\{\bm{w}\mid\|\bm{w}-\tilde{\bm{w}}\|_{2}\leq S\}$ ($S>0$).
With L2-regularization, since $\rho^{*}(\bm{v})=\frac{1}{2\lambda}\|\bm{v}\|_{2}^{2}$, the duality gap (8) is calculated as
$\sum_{i=1}^{n}w_{i}[\ell_{y_{i}}(\check{X}_{i:}\bm{\beta}^{*(\tilde{\bm{w}})})+\ell^{*}_{y_{i}}(-\alpha^{*(\tilde{\bm{w}})}_{i})]+\frac{\lambda}{2}\|\bm{\beta}^{*(\tilde{\bm{w}})}\|_{2}^{2}+\frac{1}{2\lambda}\bm{w}^{\top}(\bm{\alpha}^{*(\tilde{\bm{w}})}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})(\bm{\alpha}^{*(\tilde{\bm{w}})}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})^{\top}\bm{w}.$ (9)
Since $(\bm{\alpha}^{*(\tilde{\bm{w}})}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})(\bm{\alpha}^{*(\tilde{\bm{w}})}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})^{\top}$ is positive semidefinite, we can see that (9) is a quadratic convex function with respect to $\bm{w}$. As a method to solve a constrained maximization of (9) in ${\cal W}:=\{\bm{w}\mid\|\bm{w}-\tilde{\bm{w}}\|_{2}\leq S\}$, let us apply the method of Lagrange multipliers. Then we can prove that all stationary points can be algorithmically computed, and therefore the maximization can be also computed. In fact,
Theorem 4.1.
The following maximization can be solved algorithmically:
$\max_{\bm{w}\in{\cal W}}\bm{w}^{\top}A\bm{w}+2\bm{b}^{\top}\bm{w},$ (10)
$\text{where}\quad{\cal W}:=\{\bm{w}\in\mathbb{R}^{n}\mid\|\bm{w}-\tilde{\bm{w}}\|_{2}\leq S\},$
$\quad\tilde{\bm{w}}\in\mathbb{R}^{n},\quad\bm{b}\in\mathbb{R}^{n},\quad S>0,$
[…]
See Appendix D.2 for the proof. Although the maximization requires $O(n^{3})$ time, the result is valuable since conventional SSS methods cannot derive SSS rules for a range of $\bm{w}$. After maximizing (9) by Theorem 4.1, we can compute (7) and therefore the DRSSS rule in Theorem 3.5 can be computed.
4.2 DRSSS for L2-regularization using kernels
In this section we show the condition to apply both the kernel method and DRSSS, and show the explicit computations when L2-regularization is employed.
In this paper we omit the detailed definition of the kernel method; however, within this context, we define it as the scenario where input variables for any instance (e.g., elements of $X$) or primal model parameters (e.g., $\bm{\beta}^{*(\tilde{\bm{w}})}$) cannot be explicitly obtained, yet inner products between samples (e.g., $X_{i:}X_{i^{\prime}:}^{\top}$) are accessible. We call that weighted RERM in Section 2.1 can be kernelized if all necessary quantities can be computed even under the circumstances described above. For detailed formulation of kernels, see Chapter 2 of [20] for example. In Appendix D.1 we will discuss the condition of regularization functions so that weighted RERM can accept both kernelization and SSS (although it is not obvious to extend to DRSSS).
In this section, as an example, we show DRSSS computation when kernelized with L2-regularization. To do this, we need to compute (6) and (9) with kernelized setup. Noticing that the relationship (4) for L2-regularization is calculated as $\bm{\beta}^{*(\tilde{\bm{w}})}=\frac{1}{\lambda}(\tilde{\bm{w}}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})^{\top}\bm{\alpha}^{*(\tilde{\bm{w}})}=\frac{1}{\lambda}\check{X}^{\top}(\tilde{\bm{w}}\otimes\bm{\alpha}^{*(\tilde{\bm{w}})})$, we can compute (6) and (9) with kernelized as
$\check{X}_{i:}\bm{\beta}^{*(\tilde{\bm{w}})}\pm\|\check{X}_{i:}\|_{2}R=\frac{1}{\lambda}\dashuline{\check{X}_{i:}\check{X}^{\top}}(\tilde{\bm{w}}\otimes\bm{\alpha}^{*(\tilde{\bm{w}})})\pm\sqrt{\dashuline{\check{X}_{i:}\check{X}_{i:}^{\top}}}R,$
$P_{\bm{w}}(\bm{\beta}^{*(\tilde{\bm{w}})})-D_{\bm{w}}(\bm{\alpha}^{*(\tilde{\bm{w}})})=\sum_{i=1}^{n}w_{i}\left[\ell_{y_{i}}\left(\frac{1}{\lambda}\dashuline{\check{X}_{i:}\check{X}^{\top}}(\tilde{\bm{w}}\otimes\bm{\alpha}^{*(\bm{w})})\right)+\ell^{*}_{y_{i}}(-\alpha^{*(\tilde{\bm{w}})}_{i})\right]$
$+\frac{1}{2\lambda}(\bm{w}-\tilde{\bm{w}})^{\top}\mathrm{diag}(\bm{\alpha}^{*(\tilde{\bm{w}})})\dashuline{\check{X}\check{X}^{\top}}\mathrm{diag}(\bm{\alpha}^{*(\tilde{\bm{w}})})(\bm{w}-\tilde{\bm{w}}).$
Here, expressions with dashed lines denote that the quantities that can be computed even when kernelized, that is, without using $\bm{\beta}^{*(\tilde{\bm{w}})}$ or $X$ except for inner products between samples $X_{i:}X_{i^{\prime}:}^{\top}$. This concludes that DRSSS rules can be computed even kernelized with L2-regularization.
5 Numerical experiment
5.1 Experimental settings
We evaluate the performances of DRSSS across different values of acceptable weight changes $S$ and hyperparameters for regularization strength $\lambda$. Performance is measured by safe screening rate, representing the ratio of the number of screened samples to that of all samples. We examined DRSSS performances with SVM (Hinge loss + L2-regularization) under three setups:
  • Tabular data, linear prediction (no kernel)
  • Tabular data, radial basis function (RBF) kernel [20]
  • Image data, neural tangent kernel (NTK) (see Section 1)
In these experiments, we set the initial sample weights ($\tilde{\bm{w}}$ in Theorem 3.5) as $\tilde{\bm{w}}=\bm{1}_{n}$, and learn $\bm{\alpha}^{*(\tilde{\bm{w}})}$ by solving (2)222 Since the experiments do not employ validation or test dataset, we use all samples in the dataset for training. . If no kernel is employed, we compute $\bm{\beta}^{*(\tilde{\bm{w}})}$ from $\bm{\alpha}^{*(\tilde{\bm{w}})}$ by (4). Then, we set the range of weights as ${\cal W}:=\{\bm{w}\mid\|\bm{w}-\tilde{\bm{w}}\|_{2}\leq S\}$ (Section 4), where $S$ is specified as follows: First we assume the weight change that the weights for positive samples ($\{i\mid y_{i}=+1\}$) from $1$ to $a$, while retaining the weights for negative samples ($\{i\mid y_{i}=-1\}$) as $1$. Then, we defined $S$ as the size of weight change above; specifically, we set $S=\sqrt{n^{+}}|a-1|$ ($n^{+}$: number of positive samples in the training dataset). We vary $a$ within the range $0.95\leq a\leq 1.05$, assuming a maximum change of up to 5% per sample weight. Finally we apply DRSSS in Section 4.
Also, we demonstrate the difficulty of conventional SSS to achieve DRSSS in Section 5.2. Since conventional SSS can handle only specific weight change, we consider “simulating” the maximization computation $\max_{\bm{w}\in{\cal W}}$ in (7) (Theorem 3.5) by repeatedly choosing $\bm{w}$ from $\|\bm{w}-\tilde{\bm{w}}\|_{2}=S$ at random and regarding the largest one among chosen $\bm{w}$ as $R$. As a result, $R$ is underestimated and therefore samples that should not be screened out may be incorrectly screened out.
5.2 Safe screening rates for tabular data
[…]
(Model: Linear)
(Model: RBF kernel)
Figure 2: Ratio of screened samples by DRSSS for dataset “breast-cancer”.
First, we present safe screening rates for two SVM setups: linear prediction (no kernel) and the RBF kernel [20]. The RBF kernel is a famous kernel function for finite-dimensional vectors to achieve nonlinear predictions. Given the original dataset $Z\in\mathbb{R}^{n\times d^{\prime}}$ with $n$ samples and $d^{\prime}$ features, the RBF kernel can be computed as $X_{i:}X_{i^{\prime}:}^{\top}=\exp(-\|Z_{i:}-Z_{i^{\prime}:}\|_{2}/\zeta)$, where $\zeta>0$ is a hyperparameter. The datasets used in these experiments are detailed in Table 2. In this experiment, we adapt the regularization hyperparameter $\lambda$ based on the characteristics of the data. These details are described in Appendix E.1.
As an example, for the “breast-cancer” dataset, we show the DRSSS result in Figure 2 for setups of no kernel and RBF kernel. Results for other datasets are presented in Appendix E.3.
These plots allow us to assess the tolerance for changes in sample weights. Since $a=1$ means no weight change, the screening rates fall as $a$ becomes apart from 1. We can also see that the screening rates tend to be large for small $\lambda$ if $a$ is near to 1, while the falls of screening rates is slower for large $\lambda$. The former means that smaller $\lambda$ tends to have more unnecessary samples ($\alpha^{*(\tilde{\bm{w}})}_{i}=0$) if we do not consider the change of weights. On the other hand, the behavior of the latter is accounted for by Lemma 3.1 since larger $\lambda$ reduces $r(\bm{w},\kappa,\hat{\bm{\beta}},\hat{\bm{\alpha}})$. For instance, with $a=0.98$ (weight of each positive sample is reduced by two percent, or equivalent weight change in L2-norm), the sample screening rates are 0.82 (linear) and 0.27 (RBF kernel) for SVM with $\lambda=\mathrm{68.3}$. This implies that, even if the weights are changed in such range, a number of samples are still identified as redundant in the sense of prediction.
Then we show the result of simulated DRSSS by repeatedly applying conventional SSS. In Figure 4 we presented the number of unscreened samples for the dataset “breast-cancer”, $\lambda=6.83$ and $a=1.02$. By the proposed DRSSS, we identified that 198 samples are needed to be kept. However, by the simulation, even if we choose $\bm{w}$ for 1,024 times at random, only 71 samples are identified to be needed on average (the right end of the blue line), which demonstrates the difficulty of conventional SSS to apply to our DRSSS setup. Experimental details and results for other datasets are presented in Appendix E.2.
5.3 Safe screening rates with neural tangent kernel
Figure 3: Number of unscreened samples by DRSSS and by the simulation of repeating conventional SSS, for the dataset “breast-cancer”, linear prediction, $\lambda=6.83$ and $a=1.02$. Error bars are presented but very short; see Appendix E.2 for details.
Figure 4: Ratio of screened samples by DRSSS for dataset “CIFAR-10” using neural tangent kernel (NTK).
As noted in the beginning of the section, we experimented DRSSS with NTK by approximating the a neural network for images. First, we composed 5,000-sample binary classification dataset from “CIFAR-10” dataset [22], by choosing from classes “airplane” (2,529) and ‘’automobile” (2,471). Then, as a neural network for images, we set up a convolutional network of input width 1,024 (same as the number of pixels in an image), three layers of $3\times 3$ convolution filters, followed by a linear fully connected layer. With these setups, we trained NTK and obtained the kernel values for all pairs of samples by a Python library “Neural Tangents” [23]. Finally, similar to tabular datasets, we trained a weighted RERM with the hinge loss and the L2-regularization and applied DRSSS.
The result is presented in Figure 4. We can confirm a similar tendency to the ones discussed in Section 5.2, which demonstrates the validity of the proposed method.
6 Conclusions, limitations and future works
In this study, we propose the DRSSS method as a robust data reduction method for situations where the data distribution at deployment is unknown. This method effectively reduces data storage and model update costs in DR learning settings. Our technical contribution extends the SSS technique, originally developed for accelerating a single optimization problem, to be applicable across an entire class of optimization problems within a specific category. The current limitation of the proposed method is that the applicability of DRSSS, similar to SSS, is confined to convex optimization problems of a specific class. Therefore, applying it to complex models such as deep learning requires using kernel methods based on NTK, meaning it cannot be directly applied to deep learning models. Additionally, there may be cases where the removable data volume by DRSSS is insufficient, since DRSSS identifies unnecessary samples under a stringent condition that exactly the same model parameter should be obtained even when they are removed. It is, therefore, crucial to develop a framework that relaxes these requirements and allows for the removal of more data.
Acknowledgments and Disclosure of Funding
This work was partially supported by MEXT KAKENHI (20H00601, 24K15080), JST CREST (JPMJCR21D3 including AIP challenge program, JPMJCR22N2), JST Moonshot R&D (JPMJMS2033-05), JST AIP Acceleration Research (JPMJCR21U2), NEDO (JPNP18002, JPNP20006) and RIKEN Center for Advanced Intelligence Project.
Appendix A Proofs of general lemmas
In this appendix we show several lemmas used in other appendices.
Lemma A.1.
(Fenchel-Moreau theorem) For a convex function $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$, $f^{**}$ is equivalent to $f$ if $f$ is convex, proper (i.e., $\exists\bm{v}\in\mathbb{R}^{d}:~{}f(\bm{v})<+\infty$) and lower-semicontinuous.
Proof.
See Section 12 of [24] for example. ∎
As a special case of Lemma A.1, it holds if $f$ is finite ($\forall\bm{v}\in\mathbb{R}^{d}:~{}f(\bm{v})<+\infty$) and convex.
Lemma A.2.
For a convex function $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$,
  • $f^{*}$ is $(1/\nu)$-strongly convex if $f$ is proper and $\nu$-smooth.
  • $f^{*}$ is $(1/\kappa)$-smooth if $f$ is proper, lower-semicontinuous and $\kappa$-strongly convex.
Proof.
See Section X.4.2 of [25] for example. ∎
Corollary A.3.
Suppose that the regularization function $\rho$ in (1) is $\kappa$-strongly convex, which is required to apply SSS (Lemma 3.1). Then, from Lemma A.2, $\rho^{*}$ must be $(1/\kappa)$-smooth. This implies that, in this case, $\rho^{*}$ cannot be infinite.
Lemma A.4.
Suppose that $f:\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is a $\kappa$-strongly convex function, and let $\bm{v}^{*}={\rm argmin}_{\bm{v}\in\mathbb{R}^{d}}f(\bm{v})$ be the minimizer of $f$. Then, for any $\bm{v}\in\mathbb{R}^{d}$, we have
$\|\bm{v}-\bm{v}^{*}\|_{2}\leq\sqrt{\frac{2}{\kappa}[f(\bm{v})-f(\bm{v}^{*})]}.$
Proof.
See [11] for example. ∎
Lemma A.5.
For any vector $\bm{a},\bm{c}\in\mathbb{R}^{n}$ and $S>0$,
$\min_{\bm{v}\in\mathbb{R}^{n}:~{}\|\bm{v}-\bm{c}\|_{2}\leq S}\bm{a}^{\top}\bm{v}=\bm{a}^{\top}\bm{c}-S\|\bm{a}\|_{2},$ $\max_{\bm{v}\in\mathbb{R}^{n}:~{}\|\bm{v}-\bm{c}\|_{2}\leq S}\bm{a}^{\top}\bm{v}=\bm{a}^{\top}\bm{c}+S\|\bm{a}\|_{2}.$
Proof.
By Cauchy-Schwarz inequality,
$-\|\bm{a}\|_{2}\|\bm{v}-\bm{c}\|_{2}\leq\bm{a}^{\top}(\bm{v}-\bm{c})\leq\|\bm{a}\|_{2}\|\bm{v}-\bm{c}\|_{2}.$
Noticing that the first inequality becomes equality if $\exists\omega>0:~{}\bm{a}=-\omega(\bm{v}-\bm{c})$, while the second inequality becomes equality if $\exists\omega^{\prime}>0:~{}\bm{a}=\omega^{\prime}(\bm{v}-\bm{c})$. Moreover, since $\|\bm{v}-\bm{c}\|_{2}\leq S$,
$-S\|\bm{a}\|_{2}\leq\bm{a}^{\top}(\bm{v}-\bm{c})\leq S\|\bm{a}\|_{2}$
also holds, with the equality holds if $\|\bm{v}-\bm{c}\|_{2}=S$.
On the other hand, if we take $\bm{v}$ that satisfies both of the equality conditions of Cauchy-Schwarz inequality above, that is,
  • (for the first inequality being equality) $\bm{v}=\bm{c}-(S/\|\bm{a}\|_{2})\bm{a}$,
  • (for the second inequality being equality) $\bm{v}=\bm{c}+(S/\|\bm{a}\|_{2})\bm{a}$,
then the inequalities become equalities. This proves that $-S\|\bm{a}\|_{2}$ and $S\|\bm{a}\|_{2}$ are surely the minimum and maximum of $\bm{a}^{\top}(\bm{v}-\bm{c})$, respectively. ∎
Appendix B Proofs of Section 2
As the formulation of Fenchel’s duality theorem, we follow the one in Section 31 of [24].
Lemma B.1 (A special case of Fenchel’s duality theorem: $f,g<+\infty$).
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ and $g:\mathbb{R}^{d}\to\mathbb{R}$ be convex functions, and $A\in\mathbb{R}^{n\times d}$ be a matrix. Moreover, we define
$\bm{v}^{*}:=\min_{\bm{v}\in\mathbb{R}^{d}}[f(A\bm{v})+g(\bm{v})],$ (11)
$\bm{u}^{*}:=\max_{\bm{u}\in\mathbb{R}^{n}}[-f^{*}(-\bm{u})-g^{*}(A^{\top}\bm{u})].$ (12)
Then Fenchel’s duality theorem assures that
$f(A\bm{v}^{*})+g(\bm{v}^{*})=-f^{*}(-\bm{u}^{*})-g^{*}(A^{\top}\bm{u}^{*}),$
$-\bm{u}^{*}\in\partial f(A\bm{v}^{*}),$
$\bm{v}^{*}\in\partial g^{*}(A^{\top}\bm{u}^{*}).$
Proof.
Introducing a dummy variable $\bm{\psi}\in\mathbb{R}^{n}$ and a Lagrange multiplier $\bm{u}\in\mathbb{R}^{n}$, we have
$\min_{\bm{v}\in\mathbb{R}^{d}}[f(A\bm{v})+g(\bm{v})]=\max_{\bm{u}\in\mathbb{R}^{n}}\min_{\bm{v}\in\mathbb{R}^{d},~{}\bm{\psi}\in\mathbb{R}^{n}}[f(\bm{\psi})+g(\bm{v})-\bm{u}^{\top}(A\bm{v}-\bm{\psi})]$ (13)
$=-\min_{\bm{u}\in\mathbb{R}^{n}}\max_{\bm{v}\in\mathbb{R}^{d},~{}\bm{\psi}\in\mathbb{R}^{n}}[-f(\bm{\psi})-g(\bm{v})+\bm{u}^{\top}(A\bm{v}-\bm{\psi})]$
$=-\min_{\bm{u}\in\mathbb{R}^{n}}\max_{\bm{v}\in\mathbb{R}^{d},~{}\bm{\psi}\in\mathbb{R}^{n}}[\{(-\bm{u})^{\top}\bm{\psi}-f(\bm{\psi})\}+\{(A^{\top}\bm{u})^{\top}\bm{v}-g(\bm{v})\}]$
$=-\min_{\bm{u}\in\mathbb{R}^{n}}[f^{*}(-\bm{u})+g^{*}(A^{\top}\bm{u})]=\max_{\bm{u}\in\mathbb{R}^{n}}[-f^{*}(-\bm{u})-g^{*}(A^{\top}\bm{u})].$ (14)
Moreover, by the optimality condition of a problem with a Lagrange multiplier (13), the optima of it, denoted by $\bm{v}^{*}$, $\bm{\psi}^{*}$ and $\bm{u}^{*}$, must satisfy
$A\bm{v}^{*}=\bm{\psi}^{*},\quad A^{\top}\bm{u}^{*}\in\partial g(\bm{v}^{*}),\quad-\bm{u}^{*}\in\partial f(\bm{\psi}^{*})=\partial f(A\bm{v}^{*}).$
On the other hand, introducing a dummy variable $\bm{\phi}\in\mathbb{R}^{d}$ and a Lagrange multiplier $\bm{v}\in\mathbb{R}^{d}$ for (14), we have
$\max_{\bm{u}\in\mathbb{R}^{n}}[-f^{*}(-\bm{u})-g^{*}(A^{\top}\bm{u})]=\min_{\bm{v}\in\mathbb{R}^{d}}\max_{\bm{u}\in\mathbb{R}^{n},\bm{\phi}\in\mathbb{R}^{d}}[-f^{*}(-\bm{u})-g^{*}(\bm{\phi})-\bm{v}^{\top}(A^{\top}\bm{u}-\bm{\phi})]$ (15)
$=\min_{\bm{v}\in\mathbb{R}^{d}}\max_{\bm{u}\in\mathbb{R}^{n},\bm{\phi}\in\mathbb{R}^{d}}[\{(A\bm{v})^{\top}(-\bm{u})-f^{*}(-\bm{u})\}+\{\bm{v}^{\top}\bm{\phi}-g^{*}(\bm{\phi})\}]$
[…]
Likely above, by the optimality condition of a problem with a Lagrange multiplier (15), the optima of it, denoted by $\bm{u}^{*}$, $\bm{\phi}^{*}$ and $\bm{v}^{*}$, must satisfy
$A^{\top}\bm{u}^{*}=\bm{\phi}^{*},\quad\bm{v}^{*}\in\partial g^{*}(\bm{\phi}^{*})=\partial g^{*}(A^{\top}\bm{u}^{*}),\quad A\bm{v}^{*}\in\partial f(-\bm{u}^{*}).$
Lemma B.2 (Dual problem of weighted regularized empirical risk minimization (weighted RERM)).
For the minimization problem
$\bm{\beta}^{*(\bm{w})}:=\mathop{\rm argmin}\limits_{\bm{\beta}\in\mathbb{R}^{d}}P_{\bm{w}}(\bm{\beta}),\quad\text{where}\quad P_{\bm{w}}(\bm{\beta}):=\sum_{i=1}^{n}w_{i}\ell_{y_{i}}(\check{X}_{i:}\bm{\beta})+\rho(\bm{\beta}),$ ((1) restated)
we define the dual problem as the one obtained by applying Fenchel’s duality theorem (Lemma B.1), which is defined as
$\bm{\alpha}^{*(\bm{w})}:=\mathop{\rm argmax}\limits_{\bm{\alpha}\in\mathbb{R}^{n}}D_{\bm{w}}(\bm{\alpha}),\quad\text{where}\quad D_{\bm{w}}(\bm{\alpha}):=-\sum_{i=1}^{n}w_{i}\ell^{*}_{y_{i}}(-\alpha_{i})+\rho^{*}((\bm{w}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})^{\top}\bm{\alpha}).$ ((2) restated)
Moreover, $\bm{\beta}^{*(\bm{w})}$ and $\bm{\alpha}^{*(\bm{w})}$ must satisfy
$P_{\bm{w}}(\bm{\beta}^{*(\bm{w})})=D_{\bm{w}}(\bm{\alpha}^{*(\bm{w})}),$ ((3) restated)
$\bm{\beta}^{*(\bm{w})}\in\partial\rho^{*}((\bm{w}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})^{\top}\bm{\alpha}^{*(\bm{w})}),$ ((4) restated)
$\forall i\in[n]:\quad-\alpha^{*(\bm{w})}_{i}\in\partial\ell_{y_{i}}(\check{X}_{i:}\bm{\beta}^{*(\bm{w})}).$ ((5) restated)
Proof.
To apply Fenchel’s duality theorem, we have only to set $f$, $g$ and $A$ in Lemma B.1 as
$f(\bm{u}):=\sum_{i=1}^{n}w_{i}\ell_{y_{i}}(u_{i}),\quad g(\bm{\beta}):=\rho(\bm{\beta}),\quad A:=\check{X}.$
Here, noticing that
$f^{*}(\bm{u})=\sup_{\bm{u}^{\prime}\in\mathbb{R}^{n}}[\bm{u}^{\top}\bm{u}^{\prime}-\sum_{i=1}^{n}w_{i}\ell_{y_{i}}(u^{\prime}_{i})]=\sup_{\bm{u}^{\prime}\in\mathbb{R}^{n}}\sum_{i=1}^{n}\left[u_{i}u^{\prime}_{i}-w_{i}\ell_{y_{i}}(u^{\prime}_{i})\right]$
$=\sup_{\bm{u}^{\prime}\in\mathbb{R}^{n}}\sum_{i=1}^{n}w_{i}\left[\frac{u_{i}}{w_{i}}u^{\prime}_{i}-\ell_{y_{i}}(u^{\prime}_{i})\right]=\sum_{i=1}^{n}w_{i}\ell_{y_{i}}^{*}\left(\frac{u_{i}}{w_{i}}\right),$
from (12) we have
$-f^{*}(-\bm{u})-g^{*}(A^{\top}\bm{u})=-\sum_{i=1}^{n}w_{i}\ell_{y_{i}}^{*}\left(-\frac{u_{i}}{w_{i}}\right)-\rho^{*}(\check{X}^{\top}\bm{u}).$
Replacing $u_{i}\leftarrow w_{i}\alpha_{i}$, that is, $\bm{u}\leftarrow(\bm{w}\otimes\bm{\alpha})$, we have the dual problem (2).
The relationships between the primal and the dual problem are described as follows:
$-\bm{u}^{*}\in\partial f(A\bm{v}^{*})~{}\Rightarrow~{}-\bm{w}\otimes\bm{\alpha}^{*(\bm{w})}\in\partial f(\check{X}\bm{\beta}^{*(\bm{w})})~{}\Rightarrow~{}-w_{i}\alpha^{*(\bm{w})}_{i}\in w_{i}\partial\ell_{y_{i}}(\check{X}_{i:}\bm{\beta}^{*(\bm{w})})$
$\Rightarrow-\alpha^{*(\bm{w})}_{i}\in\partial\ell_{y_{i}}(\check{X}_{i:}\bm{\beta}^{*(\bm{w})}),$
$\bm{v}^{*}\in\partial g^{*}(A^{\top}\bm{u}^{*})~{}\Rightarrow~{}\bm{\beta}^{*(\bm{w})}\in\partial g^{*}(\check{X}^{\top}\bm{w}\otimes\bm{\alpha}^{*(\bm{w})})=\partial g^{*}((\bm{w}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})^{\top}\bm{\alpha}^{*(\bm{w})}).$
Appendix C Proofs of Section 3
C.1 Proof of Lemma 3.1
Proof.
[11]
[…]
C.2 Proof of Lemma 3.2
Proof.
Due to (5), if $\partial\ell_{y_{i}}(\check{X}_{i:}\bm{\beta}^{*(\bm{w})})=\{0\}$ is assured, then $\alpha_{i}^{*(\bm{w})}=0$ is assured. Since we do not know $\bm{\beta}^{*(\bm{w})}$ but know ${\cal B}^{*(\bm{w})}$ (Lemma 3.1), we can assure $\alpha_{i}^{*(\bm{w})}=0$ if $\bigcup_{\bm{\beta}\in{\cal B}^{*(\bm{w})}}\partial\ell_{y_{i}}(\check{X}_{i:}\bm{\beta})=\{0\}$ is assured. Noticing that $\partial\ell_{y_{i}}$ is monotonically increasing [^3] Since $\partial\ell_{y_{i}}$ is a multi-valued function, the monotonicity must be defined accordingly: we call a multi-valued function $F:\mathbb{R}\to 2^{\mathbb{R}}$ is monotonically increasing if, for any $t<t^{\prime}$, $F$ must satisfy “$\forall s\in F(t)$, $\forall s^{\prime}\in F(t^{\prime})$$s\leq s^{\prime}$”., we have
$\bigcup_{\bm{\beta}\in{\cal B}^{*(\bm{w})}}\partial\ell_{y_{i}}(\check{X}_{i:}\bm{\beta})=\{0\}\quad\Leftrightarrow\quad\bigcup_{\bm{\beta}\in{\cal B}^{*(\bm{w})}}\check{X}_{i:}\bm{\beta}\subseteq{\cal Z}[\ell_{y_{i}}]\quad\Leftrightarrow\quad[\min_{\bm{\beta}\in{\cal B}^{*(\bm{w})}}\check{X}_{i:}\bm{\beta},\max_{\bm{\beta}\in{\cal B}^{*(\bm{w})}}\check{X}_{i:}\bm{\beta}]\subseteq{\cal Z}[\ell_{y_{i}}]$
[…]
Appendix D Proofs and additional discussions of Section 4
D.1 Use of strongly-convex regularization functions other than L2-regularization
Let us discuss using regularization functions other than L2-regularization, on condition that it is $\kappa$-strongly convex to apply DRSSS (Lemma 3.1). In case of L2-regularization, as seen in (9), the term $\rho^{*}((\bm{w}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})^{\top}\bm{\alpha}^{*(\tilde{\bm{w}})})$ in the duality gap (8) becomes a quadratic function with respect to $\bm{w}$. However, this may largely differ for other regularization functions, even if it is $\kappa$-strongly convex (to apply DRSSS; Lemma 3.1).
As a famous example, let us see the case of the elastic net regularization [26]: with hyperparameters $\lambda>0$ and $\lambda^{\prime}>0$, the regularization function is defined and its convex conjugate is calculated as
$\rho(\bm{\beta}):=\frac{\lambda}{2}\|\bm{\beta}\|_{2}^{2}+\lambda^{\prime}\|\bm{\beta}\|_{1},$
$\rho^{*}(\bm{v})=\frac{1}{2\lambda}\sum_{j=1}^{d}\left[\max\{0,|v_{j}|-\lambda^{\prime}\}\right]^{2}.$
Here, we can see that the maximization of $\rho^{*}((\bm{w}{\raisebox{0.80002pt}{$\times$}\Box}\check{X})^{\top}\bm{\alpha}^{*(\tilde{\bm{w}})})=(1/2\lambda)\sum_{j=1}^{d}\left[\max\{0,|\bm{w}\otimes\bm{\alpha}^{*(\tilde{\bm{w}})}\otimes\check{X}_{:j}|-\lambda^{\prime}\}\right]^{2}$ with respect to $\bm{w}$ seems not to be easy.
Then we discuss a (sufficient) condition of regularization functions so that weighted RERM can accept both kernelization and SSS.
Lemma D.1.
In weighted RERM of Definition 2.1, if $\rho$ is described as follows, it can accept both SSS in Lemma (3.2) and kernelization in Section 4.2: $\rho$ is described by an increasing function ${\cal H}:\mathbb{R}_{\geq 0}\to\mathbb{R}$ and a positive constant $\kappa$ as $\rho(\bm{\beta})=(\kappa/2)\|\bm{\beta}\|_{2}^{2}+{\cal H}(\|\bm{\beta}\|_{2})$.
Proof.
According to the generalized representer theorem [27], weighted RERM can be kernelized if $\rho$ is described by a strictly increasing function ${\cal G}:\mathbb{R}_{\geq 0}\to\mathbb{R}$ as $\rho(\bm{\beta})={\cal G}(\|\bm{\beta}\|_{2})$.
Combining the condition to apply SSS (Lemma 3.1) that $\rho$ should be $\kappa$-strongly convex, we have the conclusion. ∎
However, we have to notive that, in order to extend it to DRSSS, the same problem as the example of the elastic net will appear.
D.2 Maximizing Linear and Convex Quadratic Functions in Hyperball Constraint
First we present Theorem 4.1 in more specific descriptions.
Theorem D.2.
The maximization problem in Theorem 4.1:
$\max_{\bm{w}\in{\cal W}}\bm{w}^{\top}A\bm{w}+2\bm{b}^{\top}\bm{w},$ ((10) restated)
$\text{where}\quad{\cal W}:=\{\bm{w}\in\mathbb{R}^{n}\mid\|\bm{w}-\tilde{\bm{w}}\|_{2}\leq S\},$
$\quad\tilde{\bm{w}}\in\mathbb{R}^{n},\quad\bm{b}\in\mathbb{R}^{n},\quad S>0,$
[…]
can be achieved by the following procedure. First, we define $Q\in\mathbb{R}^{n\times n}$ and $\Phi:=\mathrm{diag}(\phi_{1},\phi_{2},\dots,\phi_{n})$ as the eigendecomposition of $A$ such that $A=Q^{\top}\Phi Q$, $Q$ is orthogonal ($QQ^{\top}=Q^{\top}Q=I$). Also, let $\bm{\xi}:=-\Phi Q\tilde{\bm{w}}-Q\bm{b}\in\mathbb{R}^{n}$, and
${\cal T}(\nu)=\sum_{i=1}^{n}\left(\frac{\xi_{i}}{\nu-\phi_{i}}\right)^{2}.$ (16)
Then, the maximization (10) is equal to the largest value among them:
  • For each $\nu$ such that ${\cal T}(\nu)=S^{2}$ (see Lemma D.4), the value $\nu S^{2}+(\nu\tilde{\bm{w}}+\bm{b})^{\top}Q^{\top}(\Phi-\nu I)^{-1}\bm{\xi}+\bm{b}^{\top}\tilde{\bm{w}}$, and
  • For each $\nu\in\{\phi_{1},\phi_{2},\dots,\phi_{n}\}$ (duplication removed) such that “$\forall i\in[n]:~{}\phi_{i}=\nu\Rightarrow\xi_{i}=0$”, the value
    $\max_{\bm{\tau}\in\mathbb{R}^{n}}[\nu S^{2}+(\nu\tilde{\bm{w}}+\bm{b})^{\top}Q^{\top}\bm{\tau}+\bm{b}^{\top}\tilde{\bm{w}}],$
    $\text{subject to}\quad\forall i\in{\cal F}_{\nu}:\quad\tau_{i}=\frac{\xi_{i}}{\phi_{i}-\nu},$
[…]
    $\text{where}\quad{\cal U}_{\nu}:=\{i\mid i\in[n],~{}\phi_{i}=\nu\},\quad{\cal F}_{\nu}:=[n]\setminus{\cal U}_{\nu}.$
    (Note that the maximization is easily computed by Lemma A.5.)
The computation can be done in $O(n^{3})$ time.
This can be proved as follows.
Lemma D.3.
For the optimization problem (10), its stationary points are obtained as the solution of the following equations with respect to $\bm{w}$ and $\nu\in\mathbb{R}$:
$A\bm{w}+\bm{b}-\nu(\bm{w}-\tilde{\bm{w}})=\bm{0},$ (17)
$\|\bm{w}-\tilde{\bm{w}}\|_{2}=S.$ (18)
Also, when both (17) and (18) are satisfied, the function to be maximized is calculated as
$\bm{w}^{\top}A\bm{w}+2\bm{b}^{\top}\bm{w}=\nu S^{2}+(\nu\tilde{\bm{w}}+\bm{b})^{\top}(\bm{w}-\tilde{\bm{w}})+\tilde{\bm{w}}^{\top}\bm{b}.$ (19)
Proof.
First, $\bm{w}^{\top}A\bm{w}+2\bm{b}^{\top}\bm{w}$ is convex and not constant. Then we can show that (10) is optimized in $\{\bm{w}\in\mathbb{R}^{n}\mid\|\bm{w}-\tilde{\bm{w}}\|_{2}=S\}$, that is, at the surface of the hyperball ${\cal W}$ (Theorem 32.1 of [24]). This proves (18). Moreover, with the fact, we write the Lagrangian function with Lagrange multiplier $\nu\in\mathbb{R}$ as:
$L(\bm{w},\nu):=\bm{w}^{\top}A\bm{w}+2\bm{b}^{\top}\bm{w}-\nu(\|\bm{w}-\tilde{\bm{w}}\|_{2}^{2}-S^{2}).$
Then, due to the property of Lagrange multiplier, the stationary points of (10) are obtained as
$\frac{\partial L}{\partial\bm{w}}=2A\bm{w}+2\bm{b}-2\nu(\bm{w}-\tilde{\bm{w}})=0,$
$\frac{\partial L}{\partial\nu}=\|\bm{w}-\tilde{\bm{w}}\|_{2}^{2}-S^{2}=0,$
where the former derives (17).
Finally we show (19). If both (17) and (18) are satisfied,
[…]
$=\nu\bm{w}^{\top}(\bm{w}-\tilde{\bm{w}})+\bm{b}^{\top}\bm{w}$
$=\nu(\bm{w}-\tilde{\bm{w}})^{\top}(\bm{w}-\tilde{\bm{w}})+\nu\tilde{\bm{w}}^{\top}(\bm{w}-\tilde{\bm{w}})+\bm{b}^{\top}(\bm{w}-\tilde{\bm{w}})+\bm{b}^{\top}\tilde{\bm{w}}$
[…]
$=\nu S^{2}+(\nu\tilde{\bm{w}}+\bm{b})^{\top}(\bm{w}-\tilde{\bm{w}})+\bm{b}^{\top}\tilde{\bm{w}}$ ((19) restated)
Lemma D.4.
Under the same definitions as Theorem D.2, The equation ${\cal T}(\nu)=S^{2}$ can be solved by the following procedure: Let $\bm{e}:=[e_{1},e_{2},\dots,e_{N}]$ ($N\leq n$, $k\neq k^{\prime}\Rightarrow e_{k}\neq e_{k^{\prime}}$) be a sequence of indices such that
  1. 1.
    $e_{k}\in[n]$ for any $k\in[N]$,
  2. 2.
    $i\in[n]$ is included in $\bm{e}$ if and only if $\xi_{i}\neq 0$, and
  3. 3.
    $\phi_{e_{1}}\leq\phi_{e_{2}}\leq\dots\leq\phi_{e_{N}}$.
Note that, if $\phi_{e_{k}}<\phi_{e_{k+1}}$ ($k\in[N-1]$), then ${\cal T}(\nu)$ is a convex function in the interval $(\phi_{e_{k}},\phi_{e_{k+1}})$ with $\lim_{\nu\to\phi_{e_{k}}+0}=\lim_{\nu\to\phi_{e_{k+1}}-0}=+\infty$. Then, unless $N=0$, each of the following intervals contains just one solution of ${\cal T}(\nu)=S^{2}$:
  • Intervals $(-\infty,\phi_{e_{1}})$ and $(\phi_{e_{N}},+\infty)$.
  • Let $\nu^{\#(k)}:={\rm argmin}_{\phi_{e_{k}}<\nu<\phi_{e_{k+1}}}{\cal T}(\nu)$. For each $k\in[N-1]$ such that $\phi_{e_{k}}<\phi_{e_{k+1}}$,
    • intervals $(\phi_{e_{k}},\nu^{\#(k)})$ and $(\nu^{\#(k)},\phi_{e_{k+1}})$ if ${\cal T}(\nu^{\#(k)})<S^{2}$,
    • interval $[\nu^{\#(k)},\nu^{\#(k)}]$ (i.e., point) if ${\cal T}(\nu^{\#(k)})=S^{2}$.
It follows that ${\cal T}(\nu)=S^{2}$ has at most $2n$ solutions.
By Lemma D.4, in order to compute the solution of ${\cal T}(\nu)=S^{2}$, we have only to compute $\nu^{\#(k)}$ by Newton method or the like, and to compute the solution for each interval by Newton method or the like. We show an example of ${\cal T}(\nu)$ in Figure 5.
Proof of Lemma D.4.
We show the statements in the lemma that, if $\phi_{e_{k}}<\phi_{e_{k+1}}$ ($k\in[N-1]$), then ${\cal T}(\nu)$ is a convex function in the interval $(\phi_{e_{k}},\phi_{e_{k+1}})$ with $\lim_{\nu\to\phi_{e_{k}}+0}=\lim_{\nu\to\phi_{e_{k+1}}-0}=+\infty$. Then the conclusion immediately follows.
The latter statement clearly holds. The former statement is proved by directly computing the derivative.
$\frac{d}{d\nu}{\cal T}(\nu)=\frac{d}{d\nu}\sum_{i=1}^{n}\left(\frac{\xi_{i}}{\nu-\phi_{i}}\right)^{2}=-2\sum_{i=1}^{n}\frac{\xi_{i}^{2}}{(\nu-\phi_{i})^{3}}.$
It is an increasing function with respect to $\nu$, as long as $\nu$ does not match any of $\{\phi_{i}\}_{i=1}^{n}$ such that $\xi_{i}\neq 0$. So it is convex in the interval $\phi_{e_{k}}<\nu<\phi_{e_{k+1}}$. ∎
Figure 5: An example of the expression ${\cal T}(\nu)$ (black solid line) in Lemmas D.2 and D.4. Colored dash lines denote terms in the summation $(\xi_{e_{k}}/(\nu-\phi_{e_{k}}))^{2}$. We can see that, given an interval $(\phi_{e_{k}},\phi_{e_{k+1}})$ ($k\in[N-1]$), the function is convex.
Now we are prepared to prove Theorem D.2.
Proof of Theorem D.2.
The condition (17) is calculated as
$A\bm{w}+\bm{b}=\nu(\bm{w}-\tilde{\bm{w}}),$
$(A-\nu I)(\bm{w}-\tilde{\bm{w}})=-A\tilde{\bm{w}}-\bm{b}.$
Here, let us apply eigendecomposition of $A$, denoted by $A=Q^{\top}\Phi Q$, where $Q\in\mathbb{R}^{n\times n}$ is orthogonal ($QQ^{\top}=Q^{\top}Q=I$) and $\Phi:=\mathrm{diag}(\phi_{1},\phi_{2},\dots,\phi_{n})$ is a diagonal matrix consisting of eigenvalues of $A$. Such a decomposition is assured to exist since $A$ is assumed to be symmetric and positive semidefinite. Then,
$(Q^{\top}\Phi Q-\nu I)(\bm{w}-\tilde{\bm{w}})=-Q^{\top}\Phi Q\tilde{\bm{w}}-\bm{b},$
$Q^{\top}(\Phi-\nu I)Q(\bm{w}-\tilde{\bm{w}})=-Q^{\top}\Phi Q\tilde{\bm{w}}-\bm{b},$
$(\Phi-\nu I)\bm{\tau}=\bm{\xi},\quad(\text{where}\quad\bm{\tau}:=Q(\bm{w}-\tilde{\bm{w}}),\quad\bm{\xi}:=-\Phi Q\tilde{\bm{w}}-Q\bm{b}\in\mathbb{R}^{n},)$
$\forall i\in[n]:\quad(\phi_{i}-\nu)\tau_{i}=\xi_{i}.$ (20)
Note that we have to be also aware of the constraint
$S=\|\bm{\tau}\|_{2}=\sqrt{\bm{\tau}^{\top}\bm{\tau}}=\sqrt{(\bm{w}-\tilde{\bm{w}})^{\top}Q^{\top}Q(\bm{w}-\tilde{\bm{w}})}=\|\bm{w}-\tilde{\bm{w}}\|_{2}.$ (21)
Here, we consider these two cases.
  1. 1.
    First, consider the case when $(\Phi-\nu I)$ is nonsingular, that is, when $\nu$ is different from any of $\phi_{1},\phi_{2},\dots,\phi_{n}$. Then, from (21) we have
    $S^{2}=\|\bm{\tau}\|_{2}=\sum_{i=1}^{n}\tau_{i}^{2}=\sum_{i=1}^{n}\left(\frac{\xi_{i}}{\nu-\phi_{i}}\right)^{2}\quad\bigl{(}=:{\cal T}(\nu)\bigr{)}.$ (22)
    So, values of (10) for all stationary points with respect to $\bm{w}$ and $\nu$ (on condition that $(\Phi-\nu I)$ is nonsingular) can be obtained by computing (19) for each $\nu$ satisfying (22), that is,
    • for such $\nu$ computing $\bm{\tau}$ by (20), and
    • computing (19) as $\nu S^{2}+(\nu\tilde{\bm{w}}+\bm{b})^{\top}(\bm{w}-\tilde{\bm{w}})+\bm{b}^{\top}\tilde{\bm{w}}=\nu S^{2}+(\nu\tilde{\bm{w}}+\bm{b})^{\top}Q^{\top}\bm{\tau}+\bm{b}^{\top}\tilde{\bm{w}}$.
  2. 2.
    Secondly, consider the case when $(\Phi-\nu I)$ is nonsingular, that is, when $\nu$ is equal to one of $\phi_{1},\phi_{2},\dots,\phi_{n}$. First, given $\nu$, let ${\cal U}_{\nu}:=\{i\mid i\in[n],~{}\phi_{i}=\nu\}$ be the indices of $\{\phi_{i}\}_{i}$ equal to $\nu$ (this may include more than one indices), and ${\cal F}_{\nu}:=[n]\setminus{\cal U}_{\nu}$. Note that, by assumption, ${\cal U}_{\nu}$ is not empty. Then, all stationary points of (10) with respect to $\bm{w}$ and $\nu$ (on condition that $(\Phi-\nu I)$ is singular) can be found by computing the followings for each $\nu\in\{\phi_{1},\phi_{2},\dots,\phi_{n}\}$ (duplication excluded):
    • If $\xi_{i}\neq 0$ for at least one $i\in{\cal U}_{\nu}$, the equation (20) cannot hold.
    • If $\xi_{i}=0$ for all $i\in{\cal N}_{\nu}$, the equation (20) may hold. So we calculate $\bm{\tau}$ that maximizes (10) as follows:
      • Fix $\tau_{i}=\xi_{i}/(\phi_{i}-\nu)$ for $i\in{\cal F}_{\nu}$.
      • Set the constraint $\sum_{i\in{\cal U}_{\nu}}\tau_{i}^{2}=S^{2}-\sum_{i\in{\cal F}_{\nu}}\tau_{i}^{2}$ (due to (21)).
      • Maximize (10) with respect to $\{\tau_{i}\}_{i\in{\cal U}_{\nu}}$ under the constraints above. Here, by (19) we have only to calculate
        $\max_{\bm{\tau}\in\mathbb{R}^{n}}[\nu S^{2}+(\nu\tilde{\bm{w}}+\bm{b})^{\top}(\bm{w}-\tilde{\bm{w}})+\bm{b}^{\top}\tilde{\bm{w}}],$ (23)
        $\text{subject to}\quad\forall i\in{\cal F}_{\nu}:\quad\tau_{i}=\frac{\xi_{i}}{\phi_{i}-\nu},$
[…]
        which is easily computed by Lemma A.5. The value of the maximization result is equal to that of (10) on condition that $\nu$ is specified above.
      So, collecting these result and taking the largest one, the maximization (on condition that $(\Phi-\nu I)$ is singular) is completed.
Taking the maximum of the two cases, we have the maximization result of (10).
Finally we show the computational cost. All matrix multiplications in the computation process can be done in at most $O(n^{2})$ time, and we additionally need the following computations: the eigendecomposition of $A$ and the solution of ${\cal T}(\nu)=S^{2}$ using Newton method or the like (Lemma D.4). The former can be done in $O(n^{3})$ time. The latter requires Newton method computations for $O(n)$ time, and each of Newton method computation requires evaluations of ${\cal T}(\nu)$ that requires $O(n^{2})$ time. So, if we assume that the number of repetition of Newton method is constant, it can be done in $O(n^{2})$ time. This concludes that the computational cost is $O(n^{3})$. ∎
Appendix E Details of experiments
E.1 Data and Learning Setup
The criteria of selecting datasets (Table 2) and detailed setups are as follows:
  • All of the datasets are downloaded from LIBSVM dataset [21]. We used training datasets only if test datasets are provided separately (“splice”, “svmguide1” and “madelon”).
  • We selected datasets from LIBSVM dataset containing 100 to 10,000 samples, 100 or fewer features, and the area under the curve (AUC) of the receiver operating characteristic (ROC) is 0.9 or higher for the regularization strengths ($\lambda$) we examined so that they tend to facilitate more effective sample screening.
  • In the table, the column “$d$” denotes the number of features including the intercept feature (Remark 2.2).
The choice of the regularization hyperparameter $\lambda$, based on the characteristics of the data, is as follows: We set $\lambda$ as $n$, $n\times 10^{-0.5}$, $n\times 10^{-1.0}$, $\ldots$, $n\times 10^{-3.0}$. (For DRSSS with DL, we set 1000 instead of $n$.) This is because the effect of $\lambda$ gets weaker for larger $n$.
The choice of the hyperparameter in RBF kernel is fixed as follows: we set $\zeta=d^{\prime}*\mathbb{V}(Z)$ as suggested in sklearn.svm.SVC of scikit-learn [28], where $\mathbb{V}$ denotes the elementwise sample variance.
E.2 Simulation Setup
In Section 5.2, the experiment of the “simulated” maximization of $\max_{\bm{w}\in{\cal W}}$ in (7) (Theorem 3.5) is computed as follows.
  • First, we chose a vector $\bm{w}$ from $\|\bm{w}-\tilde{\bm{w}}\|_{2}=S$ (not $\|\bm{w}-\tilde{\bm{w}}\|_{2}\leq S$; see (18) in Lemma D.3) with the following procedure: first we draw $\bm{v}$ from a standard normal distribution, then we set $\bm{w}\leftarrow S\bm{v}/\|\bm{v}\|_{2}$.
  • We repeated this for specified number of trials ${\cal Q}$, which we denote by $\bm{w}^{[1]}$, $\bm{w}^{[2]}$, $\ldots$, $\bm{w}^{[{\cal Q}]}$.
  • We define the “simulated” maximization of $\max_{\bm{w}\in{\cal W}}f(\bm{w})$ ($f$: any function) as $\max_{q\in[{\cal Q}]}f(\bm{w}^{[q]})$.
In the experiment, for each experimental setup, we tried this for 10 times, and in the plot of the result we draw its average and standard deviation.
E.3 All Experimental Results of Section 5.2
First, we show all experimental results of safe sample screening rates for linear and RBF-kernel SVMs in Figure 6.
Then, we show experimental results of unscreened samples of DRSSS by the proposed method and simulations in Figure 7. Note that, because we have the same plot as 4 for each choice of $\lambda$, $a$ and the kernel function, it is difficult to present all results. So we present for only these setups: $\lambda\in\{n,0.01n\}$, $a=1.02$ and no kernel function (linear prediction).
Dataset: australian, Model: LinearRefer to caption Dataset: australian, Model: RBF kernelRefer to caption
Dataset: breast-cancer, Model: LinearRefer to caption Dataset: breast-cancer, Model: RBF kernelRefer to caption
Dataset: heart, Model: LinearRefer to caption Dataset: heart, Model: RBF kernelRefer to caption
Dataset: ionosphere, Model: LinearRefer to caption Dataset: ionosphere, Model: RBF kernelRefer to caption
Dataset: sonar, Model: LinearRefer to caption Dataset: sonar, Model: RBF kernelRefer to caption
Dataset: splice, Model: LinearRefer to caption Dataset: splice, Model: RBF kernelRefer to caption
Dataset: svmguide1, Model: LinearRefer to caption Dataset: svmguide1, Model: RBF kernelRefer to caption
Figure 6: Safe sample screening rates for linear and RBF-kernel SVMs, under the settings described in Section 5 and Appendix E.1.
Dataset: australian, $\lambda=0.01n=6.9$Refer to caption Dataset: australian, $\lambda=n=690.0$Refer to caption
Dataset: breast-cancer, $\lambda=0.01n=6.83$Refer to caption Dataset: breast-cancer, $\lambda=n=683.0$Refer to caption
Dataset: heart, $\lambda=0.01n=2.7$Refer to caption Dataset: heart, $\lambda=n=270.0$Refer to caption
Dataset: ionosphere, $\lambda=0.01n=3.51$Refer to caption Dataset: ionosphere, $\lambda=n=351.0$Refer to caption
Dataset: sonar, $\lambda=0.01n=2.08$Refer to caption Dataset: sonar, $\lambda=n=208.0$Refer to caption
Dataset: splice, $\lambda=0.01n=10.0$Refer to caption Dataset: splice, $\lambda=n=1000.0$Refer to caption
Dataset: svmguide1, $\lambda=0.01n=30.89$Refer to caption Dataset: svmguide1, $\lambda=n=3089.0$Refer to caption
Figure 7: Numbers of samples not screened out, with proposed DRSSS and the simulated ones (see Appendix E.2 for details). Error bars are standard deviations of 10 trials.